OpenStudy (anonymous):

Give the rectangular coordinates: (-6, -150°).

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} (-6&, &-150^o)\\ ({\color{brown}{ r}}&, & \quad {\color{brown}{ \theta}}) \end{array}\qquad \begin{array}{llll} x={\color{brown}{ r}}cos({\color{brown}{ \theta}}) \\ \quad \\ y={\color{brown}{ r}}sin({\color{brown}{ \theta}}) \end{array}\implies (x,y)\)

OpenStudy (goformit100):

Hello, and A Warm Welcome to Open Study!